OpenStudy (anonymous):

Actually, I stand corrected with my question. it is how does cos(-5pi/3)= -square root of 2/2?

OpenStudy (amistre64):

its proved usually by drawing the angle inside the unit circle

OpenStudy (anonymous):

i saw that unit circle and closed the trig book in fear and ran the other way!!! Do yo have to use the unit circle? If so, how does one read it? Btw, hello again from Illinois. Back for more torture, huh? ; )

OpenStudy (amistre64):

you can use a triangle as well, bu the unit scircle is a better fit; im drawing one u p for you right now

OpenStudy (amistre64):

cos -5pi/3 is just about a full circle so; we can call it pi/3 = 60 cos(60) = 1/2 so i dont see a way to 'prove' it becasue its contradictory it first glance

OpenStudy (anonymous):

yes you have to use the unit circle. put the 'reference triangle' inside it. or just use a cheat sheet. best one is here http://tutorial.math.lamar.edu/cheat_table.aspx

OpenStudy (anonymous):

actually it does not!

OpenStudy (amistre64):

the cos of that angle = 1/2 regardless of any other measures

OpenStudy (anonymous):

\[cos(-\frac{5\pi}{3})+=\frac{1}{2}\]

OpenStudy (anonymous):

so looking at this unit circle. i see on there that 225 degrees, for example, is 5pi/4. but what if i wanted cos or sin? like in my original ? which asks cos(-5pi/4). does it not matter about the cos or sin? will the answer in square root always be the same?

OpenStudy (anonymous):

\[= \]cos(-5pi/3) =cos(5pi/3) =\[cos( 2 \pi - 5\pi/3)\] \[=\cos(\pi / 3)\] = 1/2

OpenStudy (amistre64):

cos is the x value of the angle dropped down to the x axis

OpenStudy (amistre64):

sin is the y value

OpenStudy (amistre64):

like this

OpenStudy (anonymous):

hey are you going to be around today? I am goiong to go investigate this unit circle and maybe I can get somewhere with it then! Thanks again, of course! Too bad you live so far away; i would pay a ton of money for your tutoring math abilities!

OpenStudy (amistre64):

a ton eh ;) i should be around till about 5. cant say much for after that tho